### v2

@@ 4 Experiment III: Replicating Experiment I & II on the Latest o1 Model
Given OpenAI’s latest release of the o1-preview model, which focuses on enhancing logical reasoning capabilities [6], we have repeated the above experiments with o1. The results are shown in Figure 8.
@@
In terms of replicating Experiment I, changing the numbers in the problem statements did not significantly affect the accuracy of o1’s answers, which aligns with expectations. This observation suggests that (1) o1 has enhanced numerical computation capabilities compared to previous LLMs, and (2) the specific number given in the problem statement does not influence the method o1 chooses to solve the problem. This might imply that o1 has potentially incorporated prompts like “list the equations before solving” into its built-in thought process. Nevertheless, this speculation cannot be officially confirmed, as OpenAI has not released technical details regarding o1’s improved reasoning capabilities.
@@
However, when replicating Experiment II, the average accuracy rate of o1 was only 10%, which was even lower than that of most previous LLMs. After making minimal textual modifications while altering the mathematical principles underlying the original CRT1, CRT2, and CRT3 problems, o1 persisted in selecting problem-solving approaches based on the mathematical principles corresponding to the original problems. The errors stemming from this persistence constituted 100%, 68.42%, and 75.86% of the incorrect responses to the three types of CRT problems, respectively. This may indicate that technologies such as CoT, prompt engineering, and even fine-tuning tailored to specific datasets, cannot fundamentally enhance the ability of LLMs to comprehend mathematical problems. The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes (such as auto-regressive next-token predictors), leading to their thought patterns being deeply ingrained.
@@
Figure 8: Accuracy of the o1 model in Experiment I and Experiment II.
@@ 5 Conclusion
This paper draws on the classic CRT problems from human psychology to conduct an empirical study on the “emergence” of mathematical capabilities in mainstream LLMs. By constructing forward experiments (Experiment 1) and reverse experiments (Experiment 2), the conclusions were starkly different from mainstream views. Specifically:
(1) LLMs tend to match problem-solving strategies based on textual similarity rather than truly understanding the principles of mathematical problems.
(2) Even with the introduction of the CoT reasoning or improving reasoning ability through specialized training (e.g., o1), such methods cannot entirely adjust the LLMs’ inherent but incorrect problem-solving approaches.
This paper focuses on mathematical ability, which is the foundation of all scientific research, as its central point of study. It provides a counterexample to the overexpectation of the “emergence phenomenon” of LLM capabilities and the CoT method, both of which are currently receiving much attention.
@@ Appendix A Examples
Table 2: Examples of Determining Incorrect Answers for Four Types of Problems
Problem Type Incorrect Answer Example Error Type
A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover half of the yogurt? Please think step by step.”, ”answer”: ”If the colony doubles in size every day, then on the first day it covers half of the yogurt. On the second day, it doubles in size and covers the other half of the yogurt, so now the yogurt is fully covered. Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
B {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”Let’s break down the problem step by step: 1. The colony triples daily: This means that if the initial size of the colony is 1 unit, it will be 3 units the next day, 9 units the following day, 27 units the day after, and so on. 2. It takes 4 days for the colony to completely cover the yogurt: This means that the colony reaches its maximum size on the 4th day, which is 27 units. 3. We need to find out how many days it takes for the colony to cover 1/27 of the yogurt: Since the colony triples in size daily, on the third day, the colony would be 9 units, which is 1/3 of the complete coverage. So, it means that on the third day the yogurt is 1/3 covered. Therefore, it takes 3 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 13:22:38”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days. If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”First, let’s figure out how many times the colony triples in size in the 27 days. Since the colony triples every three days, it will triple in size 27/3 = 9 times. Now, we know that the colony starts at 1/27 of the yogurt and triples in size 9 times. So, to find out how many days it takes for the colony to cover 1/27 of the yogurt, we need to divide 27 by 9, which equals 3. Therefore, it takes 3 days for the colony to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 15:55:15”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days. If it takes 4Y days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? X and Y are both numbers, you can use them to represent the final answer. Please think step by step.”, ”answer”: ”To find out how many days it takes for the patch to cover 1/27 of the yogurt, we can first find the number of times the colony triples in size to cover the entire yogurt. If the colony triples in size every 3X days, then in 4Y days, the colony would have tripled in size 4Y / 3X times. To find out how many days it takes for the colony to cover 1/27 of the yogurt, we can use the concept of exponential growth. If the colony triples every 3X days, then after n periods of 3X days, the size of the colony would be $(3^{n})$ times the original size. So, in order to find out how many days it takes for the patch to cover 1/27 of the yogurt, we can set up the equation: $3^{(4Y/3X)}$ = 1/27 Now we need to solve for n, the number of periods of 3X days it takes for the colony to cover 1/27 of the yogurt. $3^{(4Y/3X)}=3^{-3}$ Since the bases are the same, we can set the exponents equal to each other. 4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps and mathematical operations
Table 3: Examples of Determining Correct Answers for Four Types of Problems

### v4

@@ 4 Experiment III: Replicating Experiment I & II on the Latest o1 Model
Given OpenAI’s latest release of the o1-preview model, which focuses on enhancing logical reasoning capabilities[23], we have repeated the above experiments with o1. The results are shown in Fig. 9.
@@
In terms of replicating Experiment I (Fig. 9A), changing the numbers in the problem statements did not significantly affect the accuracy of o1’s answers. This observation suggests that the specific number given in the problem statement does not affect the method that o1 selects for solving the problem. This might imply that o1 has potentially incorporated prompts like “list the equations before solving” into its built-in thought process. Nevertheless, this speculation cannot be officially confirmed, as OpenAI has not released technical details regarding o1’s improved reasoning capabilities.
However, when replicating Experiment II (Fig. 9B), the average accuracy rate of o1 was only 10.0%. After the mathematical principles of the CRT problem are altered while modifying the textual description as little as possible, o1 persisted in selecting problem-solving approaches based on the mathematical principles corresponding to the original problems. The errors stemming from this persistence constituted 100%, 68.4%, and 75.7% of the incorrect answers to the three types of CRT problems, respectively.
@@
As shown in the Fig. 9C, on average, 82.5% of the incorrect answers were due to treating the modified problem as the original one and providing the same answer as for the original problem. This may indicate that technologies such as CoT, prompt engineering, and even fine-tuning tailored to specific datasets (e.g. the o1 model), cannot fundamentally enhance the ability of LLMs to comprehend mathematical problems. The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes, such as the adoption of auto-regressive next-token predictors, resulting in their thought patterns being deeply ingrained, akin to human intuition (System 1) rather than logical reasoning (System 2).
@@
Figure 9: Average Accuracy of the o1 model in Experiment I (A) and Experiment II (B), and the average proportion of incorrect answers in Experiment II (C).
5 Related Works
Some previous empirical research has questioned the reasoning abilities of LLMs[26, 25, 27, 24], particularly in solving mathematical problems. In contrast, the contributions of this paper are more pronounced in the following aspects:
Most of the previous related studies were published before the release of the o1 model (e.g.[26, 27, 24]), therefore did not assess this model, which is renowned for its reasoning abilities. The latest experimental evidence provided in this paper demonstrates that similar conclusions remain valid for the newly released o1 model.
Although some previous studies have found that some LLMs exhibit overfitting and memorization effects during the reasoning process, leading to a decline in performance on modified datasets, the degree of decline was generally not significant (e.g.[27] , less than 13%), and not all LLMs experienced a decrease. Consequently, these studies were cautious in drawing conclusions, given that even for human testers, a moderate decline is inevitable. However, in this paper, the accuracy of all LLMs drops by more than 50% on average, providing more convincing experimental evidence. We also analyze the incorrect answers generated by LLMs, demonstrating that the reasons for their decline in accuracy are significantly different from those of humans.
Some previous studies have also mentioned human cognitive science to compare and analyze the thinking patterns of LLMs[26, 25]. However, their datasets used for testing mathematical reasoning ability are purely mathematical and unrelated to cognitive science itself. In contrast, our dataset itself is a mathematical test dataset from cognitive science, originally used to study the roles of human reasoning (System 2) and intuition (System 1). Therefore, this paper offers greater interpretability and inspiration from a psychological perspective.
@@ 6 Conclusion
This paper draws on the classic CRT problems from human psychology to conduct an empirical study on the “emergence” of mathematical capabilities in mainstream LLMs. It aims to test whether LLMs possess genuine mathematical reasoning capabilities and to provide more interpretable explanations for the causes from the perspective of human psychology. By constructing forward experiments (Experiment I) and reverse experiments (Experiment II), we obtained conclusions that are starkly different from mainstream views. Specifically:
  • LLMs tend to match problem-solving strategies based on textual similarity rather than truly understanding the underlying principles of mathematical problems. This process is more akin to human intuition (System 1) rather than logical reasoning (System 2).
  • Even with the introduction of CoT or specialized training to enhance reasoning abilities (e.g., o1), such methods cannot fundamentally alter the problem-solving thinking patterns of LLMs to endow them with human-like logical reasoning abilities (System 2). This is because the prevailing paradigm for training and fine-tuning LLMs (i.e., generating the next token with the highest probability) is, at its core, similar to human intuition (System 1).
This study conducts an empirical analysis of the “emergence” of LLMs’ mathematical reasoning abilities from a psychological perspective. We hope that this study can reduce overblown expectations of LLMs’ capabilities and stimulate more empirical research to objectively evaluate the current limitations of LLMs’ abilities.
@@ Appendix A Examples
Table 2: Examples of Determining Correct Answers for Four Types of Problems - I
@@
Table 3: Examples of Determining Correct Answers for Four Types of Problems - II
@@
Table 4: Examples of Determining Incorrect Answers for Four Types of Problems - I
Problem Type Incorrect Answer Example Error Type
A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover half of the yogurt? Please think step by step.”, ”answer”: ”If the colony doubles in size every day, then on the first day it covers half of the yogurt. On the second day, it doubles in size and covers the other half of the yogurt, so now the yogurt is fully covered. Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”} incorrect solution steps
B {”index”: 33, ”query”: ”An iceberg is forming and its surface triples every year. If the iceberg takes 10 years to grow to 27 square miles, how many years are required for it to grow to 1 square miles? Please think step by step.”, ”answer”: ”First, let’s find out how much the surface area of the iceberg increases each year. If the surface triples every year, then we can represent the surface area of the iceberg over the years as follows: 1 year: x 2 years: 3x 3 years: 3*3x = 9x 4 years: 3*9x = 27x From this pattern, we see that the surface area of the iceberg increases by a factor of 3 each year, so we can represent the surface area after n years as $3^{n}*x$. We are given that after 10 years, the iceberg has a surface area of 27 square miles, so we can set up the equation: $27=3^{1}0*x$ Solving for x: $x=27/3^{1}0$ Now that we know the value of x, we can use it to find out how many years it takes for the iceberg to grow to 1 square mile: $1=3^{n}*(27/3^{10})$ Solving for n: $3^{n}=1/(27/3^{10})$, $273^{n}=1*3^{10}/$ 27, $273^{n}=3^{10}/$ 27, $n=log3(3^{10}/27)$; $n\approx 5.07$ So, it takes approximately 5.07 years for the iceberg to grow to 1 square mile.”, ”time”: ”2024-08-20 18:08:33”, ”Model”: ”gpt-3.5-turbo-0125”} mathematical calculation errors
C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days. If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”Here’s how to solve this problem step-by-step: **1. Work Backwards:** * Since the colony triples in size every three days, it means it was 1/3 the size three days before it was full. * If it takes 27 days to cover the yogurt, it means it was 1/3 the size on day 24 (27 - 3). * It was 1/3 the size again on day 21 (24 - 3). * And again on day 18 (21 - 3). * And again on day 15 (18 - 3). * And again on day 12 (15 - 3). * And again on day 9 (12 - 3). * And again on day 6 (9 - 3). * And again on day 3 (6 - 3). **2. Identify the 1/27 Point:** * Since the colony triples in size, to be 1/27 of the full size, it means the colony was 1/3 * 1/3 * 1/3 = 1/27 the size three days before it was 1/3 the size. **3. Find the Answer:** * We found that the colony was 1/3 the size on day 3. * Three days before that (day 0) the colony was 1/27 the size. **Therefore, it takes 0 days for the colony to cover 1/27 of the yogurt. It was already that size at the start.** ”, ”time”: ”2024-08-13 15:56:05”, ”Model”: ”gemini-1.5-flash-latest”} incorrect solution steps
Table 5: Examples of Determining Incorrect Answers for Four Types of Problems - II
Problem Type Incorrect Answer Example Error Type
D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days. If it takes 4Y days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? X and Y are both numbers, you can use them to represent the final answer. Please think step by step.”, ”answer”: ”To find out how many days it takes for the patch to cover 1/27 of the yogurt, we can first find the number of times the colony triples in size to cover the entire yogurt. If the colony triples in size every 3X days, then in 4Y days, the colony would have tripled in size 4Y / 3X times. To find out how many days it takes for the colony to cover 1/27 of the yogurt, we can use the concept of exponential growth. If the colony triples every 3X days, then after n periods of 3X days, the size of the colony would be $(3^{n})$ times the original size. So, in order to find out how many days it takes for the patch to cover 1/27 of the yogurt, we can set up the equation: $3^{(4Y/3X)}$ = 1/27 Now we need to solve for n, the number of periods of 3X days it takes for the colony to cover 1/27 of the yogurt. $3^{(4Y/3X)}=3^{-3}$ Since the bases are the same, we can set the exponents equal to each other. 4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”} incorrect solution steps